Bill243:

A state park is designed in a circular pattern as shown. A park ranger drives along the circular path from the ranger station to the nature center and then from the nature center to the petting zoo. How far does he drive to the nearest tenth of a mile?

Mercury:



Vocaloid:

assuming the most efficient path:  The ranger covers a 90 + 40 = 130 degrees arc around the circle. The ranger traveled 130 out of a full 360 degree circle, so he travelled (130/360) of the entire circumference 2 * pi * r Therefore, apply arc = 2*pi*r * (130/360 degrees) to find the distance traveled.